OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 3 is four more than –11. What is the number? A. –45 B. –21 C. -7/3 D. 45

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

A number, let's call it n, is divided by three. \[\frac{n}{3}\] -11 plus 4 is equal to that value. \[\frac{n}{3}=-11+4\]

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} \textit{A number divided by 3 }\implies \cfrac{x}{3} \\ \quad \\ \textit{is four more than –11}\implies -11+4 \end{array}\implies \bf \cfrac{x}{3}=-11+4\)

OpenStudy (anonymous):

how did you get those fonts

OpenStudy (destinymasha):

not the point lol

OpenStudy (anonymous):

lol I get it...